OpenStudy (anonymous):

90 is what percent of 120?

OpenStudy (anonymous):

You certainly humour me :P

OpenStudy (anonymous):

-_-

hero (hero):

75%

OpenStudy (anonymous):

isnt the answer 3.75%

hero (hero):

You really want to know what SOPA is?

OpenStudy (anonymous):

No it is 75%

hero (hero):

I don't know how you got 3.75%

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Remember to set up your equation and then it will be simpler to solve

hero (hero):

30 = 25% 60 = 50% 90 = 75% 120 = 100%

OpenStudy (anonymous):

120x=90 x=90/120 x=.75 x=75%

OpenStudy (anonymous):

wat

hero (hero):

lol

OpenStudy (anonymous):

120 multiplied by unknown number will give you 90 120x=90

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then you want to isolate x to find its value

OpenStudy (katrinakaif):

As I mentioned before, "is" stands for equivalent to.

razor99 (razor99):

hey star srry i had to got to the hospita

OpenStudy (anonymous):

why? r u ok@rax]zor

OpenStudy (katrinakaif):

"What percent" means a variable "of" stands for multiplication 90 = n x 120

razor99 (razor99):

i had to go to the hospital everyday to change my removed toe bandgage

OpenStudy (anonymous):

Try doing the next one on your own and then katrina will check to see if you did it correctly :P

OpenStudy (katrinakaif):

Star, try learning them the way we're trying to introduce it to you. Makes our work, and your life easier =P

OpenStudy (katrinakaif):

Hahaha Pipa =D

OpenStudy (anonymous):

is it 10800%

razor99 (razor99):

again i have to got o the hos pital bye

OpenStudy (anonymous):

bye

OpenStudy (katrinakaif):

90 = n x 120 90/120 = n 9/12 = n ( Cancel the top and bottom 0's) Now simplify!

OpenStudy (anonymous):

75%

OpenStudy (katrinakaif):

If the expression was 90 = 120n: -You would have to first get "n" alone by dividing by 120

OpenStudy (anonymous):

is it 75%

OpenStudy (katrinakaif):

Correct! Great work

OpenStudy (anonymous):

thanks